# Towards User-Focused Research in Training Data Attribution for Human-Centered Explainable AI

### v1

@@ Abstract.
While Explainable AI (XAI) aims to make AI understandable and useful to humans, it has been criticised for relying too much on formalism and solutionism, focusing more on mathematical soundness than user needs. We propose an alternative to this bottom-up approach inspired by design thinking: the XAI research community should adopt a top-down, user-focused perspective to ensure user relevance. We illustrate this with a relatively young subfield of XAI, Training Data Attribution (TDA). With the surge in TDA research and growing competition, the field risks repeating the same patterns of solutionism. We conducted a needfinding study with a diverse group of AI practitioners to identify potential user needs related to TDA. Through interviews (N=10) and a systematic survey (N=31), we uncovered new TDA tasks that are currently largely overlooked. We invite the TDA and XAI communities to consider these novel tasks and improve the user relevance of their research outcomes.
@@ 1. Introduction
Understanding AI model behaviour is critical for using AI models in practice, especially in high-stakes domains such as medicine, law, or finance (Parliament, 2023). End-users require the understanding of model behaviours to make informed decisions, while developers need this to build reliable models with limited biases (Dwivedi et al., 2023). Explainable AI (XAI) aims to provide insights into AI systems’ functioning for humans. It has produced a variety of explanation methods that address different model architectures and data modalities (Guidotti et al., 2018; Dwivedi et al., 2023).
@@
Despite the progress, the XAI community has faced criticism for its predominant techno-centric focus on solutions as opposed to the problem of helping users understand model behaviour (Ehsan et al., 2021). Historically, XAI research progressed bottom-up: starting with methods providing qualitative observations, then structuring the field with quantitative evaluations, and eventually considering application in real-world environments including users (see Figure 1). Consequently, XAI was mainly driven by “[solutionism (seeking technical solutions) and] formalism (seeking abstract, mathematical solutions)” (Ehsan et al., 2021, p.2). An example of this bottom-up approach is the development of the subfield of feature attribution. Feature attribution explanations quantify each input feature’s contribution to the model prediction, highlighting key inputs. One of the first feature attribution methods in the context of deep learning was presented in 2013 introducing the visualisation of neural network input gradients as saliency map explanations and providing 15 qualitative examples as evidence that saliency maps explain (Simonyan et al., 2013). Several other methods followed, either building on this work and proposing changes to satisfy specific formal axioms (e.g., (Smilkov et al., 2017; Sundararajan et al., 2017)), or proposing model-agnostic methods in a quest for a one-fits-all solution (e.g., (Lundberg and Lee, 2017; Ribeiro et al., 2016)). Around 2017, XAI reached an inflection point where researchers started to re-assess the practicality of explanation methods and proposed standard evaluation protocols to measure how well explanation methods achieve their goal of explaining (Doshi-Velez and Kim, 2017; Adebayo et al., 2018; Bilodeau et al., 2024). Subsequently, recommendations from the social science view of explanations emphasise that explanations are part of a social process and shifted the focus of the field in 2019 (Miller, 2019). XAI was criticised for its strong solutionism with an increasing amount of works studying XAI as a sociotechnical system focusing on the user (Wang et al., 2019; Ehsan et al., 2021; Brennen, 2020; Rutjes et al., 2019; Kim et al., 2022, 2023; Fok and Weld, 2023; Kaur et al., 2020; Mei et al., 2023; Ehsan et al., 2024; Lakkaraju et al., 2022) and naming this subfield human-centred XAI (HCXAI) in 2020 (Ehsan and Riedl, 2020).
Schematic showing the bottom-up and the proposed top-down research approaches. There are three boxes vertically stacked. The bottom box reads “Methods: Studies on new algorithms. XAI example: Saliency maps (Simonyan et al., 2013)”. The middle box reads “Evaluations: Studies on protocols for comparing methods. XAI example: Sanity checks for saliency maps (Adebayo et al., 2017)”. The top box reads “Applications: Studies on methods in user-facing contexts. XAI example: Human-centred explainable AI (Ehsan and Riedl, 2020)”. There is an arrow pointing upwards to the left of the boxes, annotated with “Bottom-Up Approach -¿ Driven by techno-centric solutions” and an arrow pointing downwards to the right of the boxes, annotated with “Proposed Top-Down Approach -¿ Driven by user needs”.
@@
Despite the efforts to rectify the solutionism and formalism in XAI communities, we observe repeating patterns in sub-fields of XAI. For example, such patterns are emerging in the relatively young field of training data attribution explanations (TDA), where the model behaviour is explained based on the training data (Hammoudeh and Lowd, 2024). TDA was introduced to understand deep learning models in 2017 (Koh and Liang, 2017) and has produced several technical approaches to provide TDA explanations since then (Charpiat et al., 2019; Guo et al., 2021; Pruthi et al., 2020; Schioppa et al., 2022; Park et al., 2023; Bae et al., 2024). This development demonstrates a strong focus on seeking technical solutions in the TDA field similar to how feature attribution research started. Currently, TDA research is gaining momentum due to studies showing that TDA can be applied to address issues like data valuation, debiasing, memorisation and copyright in foundational models (Feldman and Zhang, 2020; Grosse et al., 2023; Choe et al., 2024; Jain et al., 2024; Zheng et al., 2023; Kwon et al., 2023; Lin et al., 2024; Wang et al., 2024b). Yet, the current TDA landscapes shows solutionism and formalism, with few user-focused studies, leading to a possible rift between TDA methods and user needs.
In this work, we propose future research directions for the TDA community that are entirely sourced and motivated by the potential users of TDA. We strongly believe that the actual use cases and user needs should inform and guide XAI research, echoing the call for HCXAI (Ehsan and Riedl, 2020). We break down our research question into three parts:
RQ1: What do users need TDA explanations for?
RQ2: What do users need from TDA explanations?
@@
RQ3: To what extent do existing TDA approaches align with user needs?
@@
We conduct a two-stage needfinding study with potential users of TDA technology to answer our research questions. Through a qualitative interview study with AI practitioners (N=10), we identify model developers as a key user group for TDA explanations as opposed to model users. We derive a design space for TDA explanations representing different TDA tasks (e.g., identifying the training samples with the largest effect on model behaviour when removed in contrast to when given a larger weight during training). We study which tasks are needed in model development in a subsequent mixed-methods survey study (N=31) with model developers. By considering the user and their usage scenarios, we identify tasks relevant to users but largely overlooked by current TDA research. We invite the research community to address these gaps and shift the focus of TDA research from solutionism to user-serving questions, i.e., a top-down approach. We advocate for adopting this approach in other areas of XAI and AI as a whole to avoid focusing on technically elegant solutions that may overlook user needs. Our contributions are:
@@
Through a needfinding study, we identify potential use cases, where model developers emerged as key users of TDA. We find that they require TDA explanations to be (1) flexible and adapt to the user’s intuitions in debugging and (2) reliable and stable. Generally, we observe a preference for training data group attribution over individual attribution.
As a result of studying the user needs, we identify tasks and research directions for user-focused TDA that are currently understudied.
@@ 2. Related Work
Our work proposes a top-down approach to XAI research, grounded in user needs, which positions it in human-centered explainable AI (HCXAI). We apply this approach to the relatively young subfield of training data attribution (TDA) to provide early insights into users and their needs. To this end, we utilise methods from needfinding. In the following, we relate our work to HCXAI, TDA and needfinding.
2.1. Human-centred explainable AI
@@
The aim of explainable AI (XAI) research is to study methods that explain the behaviour of AI models (Guidotti et al., 2018). The field has developed various explanation methods over the years to explain different types of tasks and models. Yet, explanations are only meaningful in front of an audience. This human aspect was often neglected in XAI studies (Rong et al., 2023; Nauta et al., 2023). To overcome this gap, Human-centred XAI (HCXAI) emphasises the human in the explanation process and focuses on user perspectives in XAI (Ehsan and Riedl, 2020). In HCXAI, explanations of black-box models are studied holistically as sociotechnical systems centred around the user (Ehsan et al., 2024). HCXAI explores how to design AI technology and explanations that focus on user understanding and practicality (Ehsan and Riedl, 2020). Work in this area covers multiple aspects that affect the user, e.g. human-AI-collaboration (Kim et al., 2023; Shen et al., 2023; Taesiri et al., 2022; Lakkaraju et al., 2022), design guidelines (Liao et al., 2020; Hadash et al., 2022; Wang et al., 2019), evaluation of human understanding of explanations (Rong et al., 2023; Kim et al., 2022; Hadash et al., 2022; Rutjes et al., 2019) and understanding user needs and requirements (Juneja et al., 2024; Kaur et al., 2020; Brennen, 2020; Jin et al., 2023). Our work contributes to the HCXAI field by proposing an early consideration of user needs to inform ongoing research in XAI methods. We focus particularly on TDA explanations which have not been explored widely in HCXAI before.
2.2. Training data attribution
@@
Training data attribution (TDA) explains model predictions by pointing to training data samples relevant to the model predictions (Hammoudeh and Lowd, 2024). TDA differs from the mainstream XAI approach known as feature attribution (FA) where the model predictions are attributed to the features of an input given to the model, disregarding the impact of training data. Interest in TDA has recently accelerated with the paradigm shift towards data-centric AI (DCAI) (Jin et al., 2024), where the importance of using the right training data is emphasised over other factors like model architecture and optimisation algorithms. Under the context of DCAI, TDA has been utilised to enhance data quality by cleaning faulty data labels (Teso et al., 2021) and detecting model biases (Pezeshkpour et al., 2022; Brunet et al., 2019; Wang et al., 2024a; Jain et al., 2024). The recent rise of foundational models, such as large language models (LLMs) and diffusion-based image-generation models, further contributed to the interest in TDA, as it has proven effective for studying the inner mechanisms as well as understanding the associated copyright and privacy leakage issues therein (Feldman and Zhang, 2020; Choe et al., 2024; Zheng et al., 2023; Grosse et al., 2023; Lin et al., 2024). With the accelerated growth, we observe again a general focus on technical and mathematical solutions in the TDA community (Koh and Liang, 2017; Charpiat et al., 2019; Guo et al., 2021; Schioppa et al., 2022; Pruthi et al., 2020; Park et al., 2023; Bae et al., 2024; Choe et al., 2024; Kwon et al., 2023), without a deep understanding of user needs, even in studies applying TDA to a downstream tasks (Feldman and Zhang, 2020; Jain et al., 2024; Zheng et al., 2023; Akyürek et al., 2022). Our work addresses this gap by studying practical needs and providing concrete recommendations for the field.
2.3. Needfinding
Needfinding is an exploratory and qualitative approach to user research to understand users’ needs beyond what simply asking users may tell (Patnaik and Becker, 1999). It aims at studying the true user needs, regardless of solutions, which may not be apparent to the users themselves. Needfinding stems from the field of design thinking (von Thienen et al., 2019), which is an iterative design approach driven by user needs. Needfinding is an iterative process involving users where the subsequent studies extract greater information about their problems and needs through field studies, interviews, and lab studies. In XAI, needfinding has been used to study general needs such as user expectations for XAI (Brennen, 2020) or end-user transparency needs in AI decision-support systems (Turri et al., 2024; Lin et al., 2024) by using semi-structured interviews, scenario-based study designs and qualitative analyses of case-study related text files. Needfinding has also been applied to more narrow questions, like extracting user needs for web-search explanations through surveys and semi-structured interviews (Juneja et al., 2024) or understanding explainability needs in human-AI collaborations (Kim et al., 2023). Our work extends this overall research direction by studying the user needs of TDA explanations through a two-stage study involving semi-structured interviews and a systematic survey study.
3. Training data attribution
Our work studies user needs for training data attribution (TDA) explanations. This section provides background on TDA. TDA is links specific model behaviour to the training data, treating the data as the root cause for learned behaviours (Hammoudeh and Lowd, 2024). By providing training data relevant to what the model has learnt, TDA gives insights into the model. TDA has been applied to enhancing data quality and reducing model biases, as well as answering questions around data valuation, memorisation and copyright issues in foundational models in the recent years (Choe et al., 2024; Zheng et al., 2023; Grosse et al., 2023; Lin et al., 2024), attracting unprecedented attention.
@@ 3. Training data attribution
The TDA community has focused on a particular instance of attributing a model prediction on a single test sample on a single training sample. More formally, let us consider a training sample and a test sample , where indicates the input and indicates the true answer the model is supposed to predict. We define the attribution score for and as the change in the correctness of the model prediction on , measured via the model loss , before and after removing the sample from the training procedure (Hampel, 1974):
@@
Computing directly by re-training a model without each training sample is computationally prohibitive. To identify the most influential training sample, the re-training has to be repeated as many times as the number of training samples. The main focus of the TDA community has been on efficient approximations of Equation 1. In 2017, Koh & Liang (Koh and Liang, 2017) introduced a gradient-based approximation called influence functions (IF). Since IF was still computationally prohibitive due to the need to compute and invert a massive Hessian matrix, subsequent works in 2020 and 2021 have focused on speeding up the algorithm by trading off precision and computational costs (Pruthi et al., 2020; Charpiat et al., 2019; Schioppa et al., 2022; Guo et al., 2021). After the advent of ChatGPT in late 2022 (OpenAI, 2022) and the boom in large language models in 2023 (Naveed et al., 2023), the community has developed further enhancements to apply the TDA methods on billion-scale LLMs (Bae et al., 2024; Choe et al., 2024; Grosse et al., 2023).
@@
We observe hints of solutionism and formalism in the development of TDA research, as the majority of the community effort is dedicated to addressing computational inefficiencies of the method itself, without questioning the practical relevance and user values of TDA as defined in Equation 1. This particular formulation of TDA has not gone without criticism within the community: a prior work (Nguyen et al., 2023) discussed the statistical insignificance of removing a single training sample in modern neural network training and (Ilyas et al., 2022) recognised the limitation and argued for examining the removal of multiple training samples. Some TDA works have explored possible applications that are potentially interesting to users such as fixing mislabelled data (Koh and Liang, 2017) or identifying brittle classes (Ilyas et al., 2022). However, these applications are still not driven by user needs. There is no known work in the TDA domain involving actual users to identify their needs to assess the relevance of the current TDA tasks and identify research directions grounded in user needs that the community should focus on.
4. Understanding user needs of TDA explanations through needfinding study
Our work studies user needs of training data attribution (TDA) explanations to facilitate user need-driven TDA research. To this end, we conduct a two-stage needfinding study. First, we identify use cases for TDA explanations using semi-structured interviews. The experimental setup is described in § 4.1. § 4.2 presents the interview findings. We derive a design space representing different types of information that may be needed from TDA in § 4.3. Second, we conduct a systematic survey study targeted at extracting user needs from this space (§ 4.4). IRB approval was obtained from one of the authors’ institutions.
4.1. Interview study
In the interview study, we conduct semi-structured interviews with users of machine learning (ML) applications outside of ML research. This serves as an exploratory study to gain an impression of potential usage scenarios and user needs for TDA RQ1 – What do users need TDA explanations for?

### v2

@@ Abstract.
Explainable AI (XAI) aims to make AI systems more transparent, yet many practices emphasise mathematical rigour over practical user needs. We propose an alternative to this model-centric approach by following a design thinking process for the emerging XAI field of training data attribution (TDA), which risks repeating solutionist patterns seen in other subfields. However, because TDA is in its early stages, there is a valuable opportunity to shape its direction through user-centred practices. We engage directly with machine learning developers via a needfinding interview study (N=6) and a scenario-based interactive user study (N=31) to ground explanations in real workflows. Our exploration of the TDA design space reveals novel tasks for data-centric explanations useful to developers, such as grouping training samples behind specific model behaviours or identifying undersampled data. We invite the TDA, XAI, and HCI communities to engage with these tasks to strengthen their research’s practical relevance and human impact.
@@ 1. Introduction
Explaining and understanding AI model behaviour is critical for the use and development of AI models in practice, especially in high-stakes domains such as medicine, law, or finance (Parliament, 2023). However, despite significant progress in Explainable AI (XAI) techniques (Guidotti et al., 2018; Dwivedi et al., 2023; Nauta et al., 2023), the community has faced criticism for its predominant techno-centric focus on solutions as opposed to the sociotechnical problem of helping users understand model behaviour (Ehsan et al., 2021; Williams, 2021; Wolf, 2019). We highlight that historically, XAI research has followed a bottom-up trajectory where methods and frameworks have built on each other: beginning with technical methods that are seemingly useful, then justifying practicality through quantitative evaluation frameworks, and ultimately extending to real-world applications through field studies and user studies (Figure 1).
@@
An example of this bottom-up trajectory is the development of the subfield of feature attribution. Feature attribution explanations quantify each input feature’s contribution to the model prediction, identifying the most influential inputs. One of the first feature attribution methods in the context of deep learning was presented in 2013, introducing the visualisation of neural network input gradients as saliency map explanations (Simonyan et al., 2013). The paper provides 15 qualitative examples as evidence that saliency maps explain. Several other methods followed, either proposing changes to satisfy specific formal axioms (e.g., (Smilkov et al., 2017; Sundararajan et al., 2017)) or proposing more widely applicable model-agnostic methods (e.g., (Lundberg and Lee, 2017; Ribeiro et al., 2016)). Around 2017, XAI reached an inflection point where researchers began reassessing the practicality of explanation methods and introduced standardized evaluation protocols to systematically measure their effectiveness in achieving their explanatory goals (Doshi-Velez and Kim, 2017; Adebayo et al., 2018; Bilodeau et al., 2024). Soon after, recommendations from the social science view of explanations emphasised that explanations are part of a social process, which shifted the field’s focus to explanation effectiveness in real-world decision-making (Miller, 2019). Increasing amount of works studying XAI started to focus on the users and XAI as part of a sociotechnical system (Wang et al., 2019; Ehsan et al., 2021; Brennen, 2020; Rutjes et al., 2019; Kim et al., 2022, 2023; Fok and Weld, 2023; Kaur et al., 2020; Mei et al., 2023; Ehsan et al., 2024; Lakkaraju et al., 2022), some criticising the “[solutionism (seeking technical solutions) and] formalism (seeking abstract, mathematical solutions) (Ehsan et al., 2021, p.1)” of early approaches. In 2020, this subfield was coined as human-centred XAI (HCXAI) (Ehsan and Riedl, 2020).
@@
Despite the efforts to rectify the solutionism and formalism in XAI communities, we observe repeating patterns in the relatively young field called training data attribution (TDA) explanations, where the model behaviour is explained based on the influence of the training data (Hammoudeh and Lowd, 2024) (Figure 2). TDA was first introduced to understand deep learning models in 2017 (Koh and Liang, 2017), and has produced several technical explanation approaches since then (Charpiat et al., 2019; Guo et al., 2021; Pruthi et al., 2020; Schioppa et al., 2022; Park et al., 2023; Bae et al., 2024; Ilyas and Engstrom, 2025). The field’s development demonstrates a strong focus on seeking technical solutions similar to the trajectory of feature attribution research. We argue that TDA research is at a critical juncture, gaining momentum for its ability to address key issues such as data valuation (Ghorbani and Zou, 2019; Choe et al., 2024), debiasing (Jain et al., 2024), memorisation and copyright in foundational models (Feldman and Zhang, 2020; Zheng et al., 2023; Lin et al., 2024; Akyürek et al., 2022) and understanding model behaviour and errors (Grosse et al., 2023; Wang et al., 2024). Yet, it remains dominated by solutionism and formalism, with limited user-focused studies, potentially creating a rift between TDA methods and users’ practical needs.
In this study, we bring users to the centre of our investigation on TDA as explanations, shifting the focus from a technology-centric perspective to a user-centric one. This approach enables researchers to develop practical solutions with better usability. We propose future research directions for the TDA and HCI community that are entirely sourced and motivated by potential users of TDA. To guide a systematic exploration of user needs, we focus on users who traditionally have access and agency over the training data, model developers, as TDA explains model behaviour by identifying how each training sample is relevant to a model outcome. This study thus aims at contributing to the study of HCXAI (Ehsan and Riedl, 2020) and, more specifically, human-centred model development tools (Kaur et al., 2020; Cabrera et al., 2023).
We break down our research question into three parts:
RQ1: What data-related explanation needs, both explicit and latent, do model developers have?
@@
RQ2: To what extent do existing TDA approaches align with model developers’ needs?
RQ3: What kind of information produced by TDA will be useful for model developers?
@@
We conducted a two-stage needfinding study with model developers to address our research questions. Through a qualitative interview study (N=6), we explore potential usage scenarios of TDA explanations. We note that while TDA has the potential to support many data-centric tasks for model developers, it was not yet widely known or used in practice. Based on the interview findings, we derive a design space for TDA explanations representing different TDA tasks (e.g., identifying the training samples with the largest effect on model behaviour when removed in contrast to when given a larger weight during training). In a subsequent scenario-based interactive user study (N=31), we study which specific data-centric tasks are needed to support testing users’ hypotheses. We identify tasks relevant to potential TDA users, but are largely overlooked by current TDA research. We invite the TDA and broader XAI research communities to address these gaps and reorient TDA research away from solutionism toward user-driven inquiry – adopting a top-down, user-centred perspective. We advocate for adopting this approach in other areas of XAI and AI as a whole to avoid focusing on technically elegant solutions that may overlook user needs.
To summarise, our work makes the following contributions:
@@
From the use cases we found, we identified research directions for user-focused TDA that are currently understudied and invite the community to study these (§ 5.2).
We outline our approach into a high-level framework for needs-based research that connects user-oriented inquiry with method development in the early phases of technology research (§ 5.3).
@@ 2. Related Work
Our work takes a user-centered approach to XAI research, building from empirical insights on training data attribution (TDA) use cases, which aligns it with human-centered explainable AI (HCXAI). To this end, we draw on needfinding methods, a user-centered research approach that helps uncover latent needs and usage scenarios, to inform the design and development of TDA tools. The following sections connect our work to existing literature in HCXAI and TDA.
2.1. HCXAI for Model Development
@@
The aim of explainable AI (XAI) research is to study methods that explain the decisions and behaviour of AI models commonly referred to as “black boxes” (Guidotti et al., 2018). The field has developed various explanation methods over the years to explain different types of tasks and models. Yet, AI explanations are only meaningful if they are interpretable by the intended users in a way that informs or supports their decision-making. This human aspect was often neglected in early XAI studies (Rong et al., 2023; Nauta et al., 2023). In the Human-centred XAI (HCXAI) field, explanations of black-box models are studied holistically as sociotechnical systems centred around the human who is interested in the explanation (Ehsan et al., 2024). HCXAI explores how to design AI technology and explanations that focus on user understanding and practicality (Ehsan and Riedl, 2020). Work in this area covers multiple aspects that affect the user, e.g. human-AI-collaboration (Kim et al., 2023; Shen et al., 2023; Taesiri et al., 2022; Lakkaraju et al., 2022), design guidelines (Liao et al., 2020; Hadash et al., 2022; Wang et al., 2019), evaluation of human understanding of explanations (Rong et al., 2023; Kim et al., 2022; Hadash et al., 2022; Rutjes et al., 2019) and understanding user needs and requirements (Juneja et al., 2024; Kaur et al., 2020; Brennen, 2020; Jin et al., 2023).
However, much XAI work has focused on system end-users who are affected by an AI system (Juneja et al., 2024; Jin et al., 2023; Kim et al., 2023; Shen et al., 2023; Taesiri et al., 2022; Lakkaraju et al., 2022) and rather few explored the need for model developer who can modify the model architecture, parameters, or training process (Kaur et al., 2020; Hong et al., 2020). Our work contributes to the HCXAI field by proposing an early consideration of user needs, especially the model developers’ needs, to inform ongoing research in XAI methods. We focus particularly on TDA explanations for model developers because they have agency over the data, which have not been explored widely in HCXAI before.
2.2. TDA as Explanation
@@
Training data attribution (TDA) explains model predictions by pointing to training data samples relevant to the model predictions (Hammoudeh and Lowd, 2024). TDA differs from the mainstream XAI approach known as feature attribution (FA) where the model predictions are attributed to the features of an input given to the model, disregarding the impact of training data. Interest in TDA has recently accelerated with the paradigm shift towards data-centric AI (DCAI) (Jin et al., 2024), where the importance of using the right training data is emphasised over other factors like model architecture and optimisation algorithms (Sambasivan et al., 2021). Under the context of DCAI, TDA has been used to enhance data quality by cleaning faulty data labels (Teso et al., 2021) and detecting model biases (Pezeshkpour et al., 2022; Brunet et al., 2019; Wang et al., 2024; Jain et al., 2024), as well as answering questions around data valuation or memorisation and copyright issues in foundational models in recent years (Choe et al., 2024; Zheng et al., 2023; Grosse et al., 2023; Lin et al., 2024). As such, TDA offers the potential of providing actionable explanations for users with dataset access, like model developers. The recent rise of foundational models, such as large language models (LLMs) and diffusion-based image-generation models, further contributed to the interest in TDA, as it has proven effective for studying the inner mechanisms as well as understanding the associated copyright and privacy leakage issues therein (Feldman and Zhang, 2020; Choe et al., 2024; Zheng et al., 2023; Grosse et al., 2023; Lin et al., 2024). With the accelerated growth, we observe again a general focus on technical and mathematical solutions in the TDA community (Koh and Liang, 2017; Charpiat et al., 2019; Guo et al., 2021; Schioppa et al., 2022; Pruthi et al., 2020; Park et al., 2023; Bae et al., 2024; Choe et al., 2024; Kwon et al., 2023; Ilyas et al., 2022; Wang et al., 2025a, b; Ilyas and Engstrom, 2025), without a deep understanding of user needs, even in studies applying TDA to a downstream tasks (Feldman and Zhang, 2020; Jain et al., 2024; Zheng et al., 2023; Akyürek et al., 2022; Li et al., 2024). Our work addresses this gap by studying practical needs and providing concrete recommendations for the field.
3. Background: What is TDA?
This section provides background on training data attribution (TDA), describing what it is and why it is an interesting explanation approach. In theory, TDA links specific model behaviour to the training data, treating the data as the root cause of learned behaviours (Hammoudeh and Lowd, 2024). TDA offers insights into the model by identifying the training data most influential to what the model has learned.
@@ 3. Background: What is TDA?
Let us consider a training sample and a test sample , where indicates the input and indicates the true answer the model is supposed to predict. An attribution method assigns a score to a training data point , as the change in the correctness of the model prediction on , indicating its importance measured via the model loss , before and after removing the sample from the training procedure (Hampel, 1974):
@@
Much of the research in the TDA community has centred on efficiently approximating Equation 1 because computing directly by re-training a model without each training sample is computationally prohibitive. To identify the most influential training sample, the re-training has to be repeated as often as the number of training samples. In 2017, Koh & Liang (Koh and Liang, 2017) introduced a gradient-based approximation called influence functions (IF) for deep models. Since IF was still computationally prohibitive due to the need to compute and invert a massive Hessian matrix, subsequent works in 2020 and 2021 have focused on speeding up the algorithm by trading off precision and computational costs (Pruthi et al., 2020; Charpiat et al., 2019; Schioppa et al., 2022; Guo et al., 2021). After the advent of ChatGPT in late 2022 (OpenAI, 2022) and the boom in large language models in 2023 (Naveed et al., 2023), the community has developed further enhancements to apply the TDA methods on billion-scale LLMs (Bae et al., 2024; Choe et al., 2024; Grosse et al., 2023; Park et al., 2023). For a survey on TDA methods, we refer the reader to Hammoudeh and Lowd (2024).
@@
We observe hints of solutionism and formalism in the development of TDA research, as the majority of the community effort is dedicated to addressing computational inefficiencies of the method itself, without questioning the practical relevance and user values of TDA as defined in Equation 1. This particular formulation of TDA has not gone without criticism within the community: a prior work (Nguyen et al., 2023) discussed the statistical insignificance of removing a single training sample in modern neural network training which affects the fragility of approximation methods (Basu et al., 2021; Bae et al., 2022) and (Ilyas et al., 2022) recognised the limitation and argued for examining the removal of multiple training samples. Some TDA works have explored possible applications that are potentially interesting to developers such as fixing mislabelled data (Koh and Liang, 2017) or identifying brittle classes (Ilyas et al., 2022). However, these applications are not driven by user needs and rather targeted toward the evaluation of TDA methods, potentially missing application scenarios and not addressing user needs.
Potential.
TDA provides data-centric information that users can make use of to gain a deeper understanding of their training datasets. Such information may be particularly interesting if the user has agency over the training dataset, meaning that they are in a position to adjust it. Hence, TDA can be especially useful to model developers when a high-quality training dataset is of crucial importance (Sambasivan et al., 2021). Developers often make sense of their models through manually inspecting training data (Cabrera et al., 2023). There are tools that enable model developers to explore datasets by offering a visual interface (Hohman et al., 2020; Piorkowski et al., 2023), but they often do not inform about the connection between model output and training data. As a data-centric approach to gaining such insight, TDA could support and enhance model developers’ data-centric work. Currently, there is no known work in the TDA domain involving actual users to identify their needs to assess the relevance of the current TDA tasks and identify research directions grounded in user needs that the community should focus on.
4. Needfinding: Asking Model Developers What They Really Need
Our work investigates the user needs surrounding training data attribution (TDA) explanations to support a user-centered development for machine learning (ML) developers and practitioners. To this end, we conducted a two-stage needfinding study: first, we ran an interview study with ML practitioners (N=6) to identify the design space of TDA explanations that represent useful data-centric information for model development. Second, we build on this space through a scenario-based interactive user study (N=31) aimed at eliciting concrete user needs and expectations. IRB approval was obtained from one of the authors’ institutions.
4.1. Interview Study Procedure
We conducted semi-structured interviews with machine learning (ML) developers who work outside academic research. The goal of the interview was to explore developers’ current and anticipated practices around training data to surface latent needs that may inform the design of TDA explanation tools.